a ball having a momentum p hits a bat and its momentum becomes -p .what is the change in the momentum of the ball?
1
Expert's answer
2013-05-16T10:07:03-0400
The change momentum is difference between initial and finalmomentum and so is equal to



p - (-p) = p+p = 2p